OpenStudy (anonymous):

if 3x^2-2x+7=0 then (x-1/3)^2= A. 20/9 B. 7/9 C. -7/9 D. -8/9 E. -20/9

OpenStudy (jdoe0001):

hmm I don't see the relation what section is this covering?

OpenStudy (jdoe0001):

almost sounds like saying if potatoes are round then is going to rain on thailand which are two unrelated elements

OpenStudy (anonymous):

true I see it that way too xD I dont know the section thats why I post it here

OpenStudy (anonymous):

go eat a chapstick

OpenStudy (anonymous):

@PleaseHelp0204 who are you saying thast and why? xd

OpenStudy (anonymous):

well i just ate 4 and now i dont know what im doing

OpenStudy (anonymous):

i feel dizzy

OpenStudy (anonymous):

hahahahaha xDDD I eat weird stuff when im bored but not to that point xD

OpenStudy (johnweldon1993):

Lol i did enjoy that @jdoe0001 So basically what i see, is a polynomial *that has roots i would assume we need to solve for* And what we get from that result....we plug those results into the 2nd expression given for 'x' and solve that

OpenStudy (anonymous):

If 3x2 - 2x + 7 = 0 then x2-(2/3)x+7/3=0 In solving by completion of the square we have (x-1/3)2=x2-(2/3)x+1/9 The equation is x2-(2/3)x++21/9=0 and so (x-1/3)2+20/9=0 and so (x-1/3)2=-20/9. No need to calculate the roots at all.

OpenStudy (anonymous):

If its not asking for anything? xd

OpenStudy (johnweldon1993):

True, i like the roots though XD But yes correct answer

OpenStudy (anonymous):

I dont get anything xd

OpenStudy (anonymous):

now that i am looking at what i did. i dont get it either lol... must be the chapsticks

OpenStudy (anonymous):

hahahahahaha xDDDD

OpenStudy (anonymous):

help john

OpenStudy (anonymous):

i have no idea

OpenStudy (johnweldon1993):

Lol oh no no you work through the pain of those chapsticks and you explain!! XD

OpenStudy (anonymous):

all right well kfsv do you know completion of the squares?

OpenStudy (anonymous):

i think you can also use the quadratic formula

OpenStudy (anonymous):

Quadratic formula would probably be easiest. (2 + / - sqrt (4 - 84)) / 6 (2 + / - sqrt (-80)) / 6 (2 + / - 4i(sqrt5) / 6 x=(1 + 2i(sqrt5)) / 3 OR (1 - 2i(sqrt5)) / 3 ((1 + 2i(sqrt5)) / 3 - 1/3)^2 = (((2i(sqrt5))/3)^2) = -20/9 ((1 - 2i(sqrt5))/3 - 1/3)^2 = (((-2i(sqrt5))/3)^2) = -20/9 -20/9

OpenStudy (anonymous):

john u think thats better

OpenStudy (anonymous):

thats too complicated D: